Large Binocular Telescope, June 2011, photo by David Allen Harvey -- in his memory, September 2012

Observatory Status

July 2012

● The SAN/NAS data storage system was installed in Tucson and on the mountain.

●LUCI1 was remounted on the telescope with a replacement detector from LUCI2, and science operations with LUCI1 resumed in March 2012.

●On-sky commissioning of the SX Adaptive Optics System was begun, and it was found to have excellent performance equivalent to or better than the DX Adaptive Optics System.

●LBTI used both sides of the telescope with the adaptive loops closed to achieve fringes (fringed diffraction-limited PSFs) on both the NOMIC and LMIRCam cameras.

●Replacement of the 20 outer wheel bearings on the enclosure rotation bogies was completed.

●The new UPS was installed to support the computers in the control room on Level 2.

●Preparation of the pier for the arrival of the PEPSI spectrograph was completed, including a new steel floor in the pier at Level 3L.

January 2012

●The second adaptive secondary has been mounted on the left side of the telescope in place of the rigid secondary.

●Tracking of non-sidereal targets was implemented in TCS Build BP13.

●Failure of the LUCI1 detector array in October 2011 caused major rearrangement of the science and technical observing schedule.

●The Adaptive Optics Science Demonstration Time (AOSDT) observing was carried out in November 2011 with the DX Adaptive Secondary, AO system #1 and the PISCES camera. The run was hampered by weather, but 1.5 nights produced some spectacular diffraction-limited images.

July 2011

●Synchronous binocular presets succeeded in acquiring guide stars at both Bent Gregorian foci using TCS Build BP9.

●The telescope systems for the most part survived unprecedented cold temperatures of -27.5 Celsius in February 2011.

●LMIRCam was mounted on LBTI for the first time and used for a series of test observations at 3 and 5 microns. The LBTI adaptive optics wavefront sensor was used with the AO secondary to provide images with Strehl ratios up to ~85%.

●The infrared camera PISCES was installed at the Right Front Bent Gregorian focus for a campaign of adaptive optics commissioning in June 2011.

January 2011

●The MODS1 optical/UV spectrometer is on the telescope and it being commissioned on-sky.

●New CISCO Nexus swithgear was installed for the mountain network.

●In October 2010, LBTI brought together light from the two sides of the telescope and produced interference fringes at wavelengths of 5, 10 & 12 microns.

●The Adaptive Optics team improved further on the spectacular image quality achieved in the spring. An H-band Strehl ratio of 90% was achieved over 20 seconds in average seeing.

●Range-balanced collimation models for LBC were tested on-sky and found to make a significant improvement to the open-loop collimation.

●A failed wheel bearing on the rotating enclosure caused a loss of 15 nights of technical observing time.

July 2010

●Adaptive Optics commissioning started work on-sky on the night of 26-May-2010 UT. The first of two adaptive secondary units is mounted on the right side of the telescope. It was very impressive that the adaptive loop closed successfully on the first attempt. By the end of the month, the AO team from Arcetri was achieving Strehl ratios up to 80% in H-band on bright reference stars. Many years of development and testing effort seem to be paying rewards of excellent performance on-sky. After only 1-week of on-sky commissioning, the system is delivering images near the theoretical expectation.

● MODS1 (Multi-Object Double Spectrograph, for the optical and near-UV) arrived safely on the mountain in mid-May, and started re-assembly and alignment in the Mountain Instrument Lab.

January 2010

● Heavy snow (1.3 m in a single storm) slowed mountain operations during January.

● The six modified hardpoints were installed in the DX primary mirror cell. These will improve the resolution and repeatability of the primary mirror positioning.

● The first adaptive optics system passed its laboratory acceptance test in December 2009, and is being packed and shipped to Arizona.

● LUCIFER Science Demonstration Time observations were carried out in December 2009. LUCIFER will now join the two prime focus cameras in the stable of science instruments being used two weeks per month.

July 2009

● The first H-band images were collected with the Adaptive Optics loop closed in the Arcetri solar tower. The setup there includes the adaptive secondary, the W-unit wavefront sensor and the infrared test camera.

● MODS 1 the optical spectrograph and imager saw first light in the laboratory in Columbus.

● Adaptive Secondary thin shell #5 was completed at the Steward Observatory Mirror Lab. This is the spare shell for the two Adaptive Secondary units.

● LUCIFER 1 commissioning on the telescope will continue after summer shutdown. The imaging mode is operational. More tests are needed in the fall for longslit and multislit spectroscopy.

● Science observing continues 2 weeks per month (except July/August) with the LBC prime focus camera while all of these other instrument developments are underway.

July 2008

● The Left Front Bent Gregorian focal station is commissioned with the rigid secondary mirror and ready for the arrival of LUCIFER 1.

● LUCIFER 1, a near-infrared imager and multi-slit spectrometer was packed and shipped from Heidelberg to Safford.

● Adaptive Optics unit #1 made the first laser fringes in the solar test tower setup in Arcetri.

● LBTI was installed on the telescope and had a successful campaign of monitoring internal fringe stability. It was only a test mounting.

● The electronics for both LBC prime focus cameras were repaired/upgraded.

July 2007

● Both 8.4m primary mirrors are installed, active control verified, coated, and freshly washed.

● Both prime focus wide-field correctors are installed with near-final alignment.

● The Blue Large Binocular Camera on the SX (left) prime focus has provided six runs of shared-risk observing time for partner astronomers. (See links to LBC, Publications.)

● The Telescope is shut down for July monsoons, enabling installation and commissioning work on the SX tertiary swing arm, the hydrostatic bearing system, and the dynamic balancing system.

● Electro-mechanical tests are underway at Microgate in Italy on the fully integrated Adaptive Optics Secondary Unit#1.